A calorimeter contains 26.0 mL of water at 13.5 ∘C . When 2.50 g of X (a substance with a molar mass of 51.0 g/mol ) is added, it dissolves via the reaction X(s)+H2O(l)→X(aq) and the temperature of the solution increases to 30.0 ∘C . Calculate the enthalpy change, ΔH, for this reaction per mole of X. Assume that the specific heat of the resulting solution is equal to that of water [4.18 J/(g⋅∘C)], that density of water is 1.00 g/mL, and that no heat is lost to the calorimeter itself, nor to the surroundings.
Answer #1


X(s) + H2O(l) ----> X(aq)

heat gained by water(q) = m*s*DT

m = mass of reaction mixture = 26+2.5 = 28.5 g

s = specific heat of water = 4.184 j/g.c

DT = 30-13.5 = 16.5

q = 28.5*4.184*16.5 = 1967.53 joule

= 1.97 kj

No of mol of X = 2.5/51 = 0.05 mole

DHrxn = -q/n

      = -1.97/0.05

      = -39.4 kj/mol